# NüshuRescue: Revitalization of the endangered Nüshu Language with AI

### v2

@@ 4.2.2 Annotation Pipeline Setup
The pipeline of our LLM-in-the-loop data annotation approach is illustrated in Figure 7, where the LLM agent is queried to translate Chinese sentences into Nüshu using few-shot prompting. In the initial round (R1), seed translation examples are sampled from the NCGold dataset, while subsequent example sets are drawn from a combination of NCGold and NCSilver after the first round of translations. Beginning in the second round (R2), newly generated samples from the previous round are shuffled, with 5 samples selected to replace the top 5 samples in the original 35-sample list, and the bottom 5 samples removed. For instance, in R2, the 35 training samples consisted of the top 5 new translations from R1, followed by 30 samples from the original set. By the third round (R3), the 35 samples included 5 new translations from R2, 5 from R1, and 25 from the original set. The control test consistently used the original 35 samples across all rounds. To ensure that each Nüshu character maps to a single Chinese character, a rule-based translation length validator is integrated into the pipeline. This validator ensures that the original and translated contents have the same length. If a length mismatch occurs, the generation is rejected, and the LLM is prompted again to resolve the issue. The model is allowed up to 7 retries before moving on and classifying the instance as a failed translation.
Figure 9: Translation of unseen characters (boxed in pink)
@@ 4.2.3 Data Annotation Outcomes
Using only 35 sentences from NCGold as seed examples, the GPT-4-Turbo model achieved an accuracy of 48.69% in perfectly translating the 50 left-out NCGold examples on a strict character-by-character match. A manual review of the translations revealed that minor translation errors were predominantly due to ambiguity in Nüshu character selection, as each Nüshu character can correspond to multiple Chinese characters, and vice versa. This complexity in mapping contributed to most of the discrepancies observed. For out-of-domain Chinese sentences, the GPT-4-Turbo model successfully translated 70.00% of sentences with lengths below 16 characters. However, it frequently refused to respond when the sentence lengths exceeded 31 characters. Figure 8 displays the number of successfully translated sentences and failed translations for each round, along with the results from the control group, which is always provided the original 35 learning samples. Figure 10 exemplifies the correct, incorrect, and refusal-to-answer responses from ChatGPT in our experiments.
@@
Table 2: Evaluation results (averaged) for our Seq2Seq MT Model, evaluated on a fixed 100 test sample
@@ 5 Applications
Leveraging the NCGold and NCSilver datasets, we present both a Nüshu-Chinese language model and a Nüshu-to-Chinese translation model to demonstrate the potential of our data-annotation framework in advancing future research on Nüshu. Our experiments also underscore the critical need for further expanding the Nüshu corpus, which will enable researchers to gain deeper insights into the language and the rich cultural heritage it embodies.
@@ 5.1 Language Modeling in Nüshu
Having obtained a set of 598 Nüshu-Chinese translation sentence pairs, we employed the FastText library Athiwaratkun et al. (2018) to train a bilingual skip-gram language model in both Nüshu and Chinese. Specifically, we separated the Nüshu-Chinese translation pairs from NCGold and NCSilver into individual sentences, shuffled them, and utilized this bilingual corpus to train the FastText model. For the FastText model, we set the embedding vector size to 300, a window size of 10, and a minimum count of 5. We used the skip-gram model (sg=1), which is well-suited for smaller datasets and rare words, as it predicts surrounding words given a target word—particularly useful for the infrequent characters in the Nüshu corpus. The model was trained for 20 epochs to ensure convergence without over-fitting, with a negative sampling rate of 10 to speed up training by focusing on relevant examples. The default learning rate of 0.05 is used for training our model.
@@ 5.2 Nüshu-to-Chinese Translation
We initially trained a Seq2Seq model for Nüshu-Chinese translation using training samples from NCGold, with a batch size of 3 and for 15 epochs. For the test set, we first shuffled the 500 sentence NCGold, and fixed the first 100 samples as the test set. Despite the small dataset size, the model’s performance was promising, comparable to models developed for other truly low-resource languages Baruah et al. (2021). Encouraged by these results, we combined NCGold with NCSilver, resulting in a total of 598 sentences, to further improve translation quality. While ChatGPT demonstrated potential in translating Chinese text to Nüshu for NCSilver, it remains a closed-source model with limited interpretability, and its outputs are often noisy, requiring multiple queries for acceptable translations. To mitigate these issues and enhance both translation accuracy and model transparency, we retrained the Seq2Seq model using the combined NCGold and NCSilver datasets. We provide further details on this process in Section 6.
@@ 5.3 Nüshu Translates to Other Languages
Given the limited availability of official Nüshu-Chinese translations, data for translating Nüshu into other languages is even more scarce. As mentioned in Section 2, Chiu (2012) published a collection of works featuring Nüshu calligraphy alongside English translations of their corresponding Chinese poems. However, this is not a direct Nüshu-to-English translation and lacks certain linguistic nuances or fidelity Rosa et al. (2017). Nevertheless, it introduces the right concept—using Nüshu as a bridge for multilingual translations. Our framework enables the construction of larger Nüshu datasets than ever before, making it possible to scale up Nüshu-to-other-language translation datasets when combined with such works. This approach opens up new avenues for expanding Nüshu’s reach across languages and enhancing the diversity of translation datasets.
@@ 6 Discussion
However, we observed that performance peaked at a train dataset of 487 sentences. This dataset includes original 400 sentences from our book translations and 87 sentences generated through Round 5 (R5). Interestingly, when data from Round 6 (R6) was added, performance actually declined. We hypothesize that this decline is due to the sentences in R6 being significantly longer than those in the rest of the dataset. As shown in Table 1, the average sentence length for R6 is 31.73, compared to 23.1 for R5, while the maximum sentence length in the original 500-sentence dataset is only 18.
Our hypothesis is that the model struggles with excessively long sentences because the training data it was exposed to primarily consisted of much shorter sentences. Seq2Seq models, particularly those using a fixed-length encoder-decoder architecture, tend to perform well on sentence structures and lengths similar to those seen during training. When presented with longer, more complex sentences, the model may struggle to effectively encode the entire sentence into a fixed representation, leading to loss of important information, truncation, or difficulty in decoding the translation. Furthermore, attention mechanisms can become less effective with longer sequences, especially when the model has not been fine-tuned or adequately trained on sentences of that length. Consequently, incorporating longer sentences without balancing the model’s exposure to a variety of sentence lengths likely contributed to the performance drop observed after R6.
@@ 7 Conclusion and Future Work
In this paper, we introduced NüshuRescue, a novel approach leveraging LLMs to expand text corpora in endangered languages, with Nüshu as a case study. Our method facilitates research on these languages, while minimizing the need for extensive human annotations. Our experiments demonstrate that NüshuRescue can generate high-quality translations from Chinese to Nüshu using only 35 manually translated sentences as the seed, demonstrating its effectiveness in producing accurate translations with minimal input. The model also generalizes effectively across different sentence lengths and text domains. Notably, NüshuRescue does not require the underlying LLM to have prior knowledge of the target language, making it adaptable to other low-resource languages. In addition, we digitized 500 Chinese-Nüshu sentence pairs from ancient Chinese poetry, which were used to train both a FastText language model and a Seq2Seq Nüshu-to-Chinese MT model, in combination with the LLM-generated translations.
This work establishes a foundation for future Nüshu-centric research. Looking ahead, NüshuRescue could be extended to other under-resourced and lesser-studied languages, enabling the expansion of corpora and supporting the development of larger, more sophisticated language models.
@@
In this paper, we introduced NüshuRescue, a LLM-based framework designed to scale up text corpora in resource-scarce languages. While our experiments demonstrated its effectiveness and generalizability, we encountered limitations in significantly expanding the Nüshu corpus due to the restricted scope of the current Nüshu-Chinese dictionary. The only available expert-validated source of Nüshu content comes from "A Compendium of Chinese Nüshu" Zhao (1992), which contains only 374 Nüshu characters and 1453 corresponding Chinese characters.
Although NüshuRescue demonstrated a strong ability to generalize to out-of-vocabulary characters, the absence of expert validation prevents us from safely releasing the generated data. Consequently, we sampled only 180 sentences from the UD Chinese GSDSimp dataset Wang et al. (2021) dataset to construct the NCSilver dataset, ensuring that all included Chinese characters were within the existing Nüshu vocabulary. However, as future studies of Nüshu incorporate more characters, NüshuRescue can be easily adapted to further expand the Nüshu corpus.
@@
In this paper, we exclusively utilized publicly available resources, including the "A Compendium of Chinese Nüshu" book Zhao (1992) and the UD Chinese GSDSimp dataset Wang et al. (2021) for our experiments. We relied on off-the-shelf LLMs, such as GPT-4-Turbo, as the backbone model for NüshuRescue. Furthermore, we have thoroughly documented the prompts used to generate Nüshu translations from the LLMs, detailing how these prompts were iteratively refined to improve the quality of the outputs.
Given these considerations, we do not foresee any ethical concerns arising from our work. A sample of the NCGold and NCSilver datasets has been made available to reviewers as part of the supplementary materials, as well as through an anonymous GitHub (https://github.com/Nushu-Project/Nushu). Upon acceptance, we will publicly release the full datasets and code to the NLP community, ensuring transparency and fostering further research in this area.
@@ Appendix A Appendix Figures
Figure A1: An example from the book (a public domain page scan included in the product description on Amazon, and therefore not subject to copyright restrictions)

### v4

@@ 4.2.2 Generation Pipeline Setup
The pipeline of our LLM-in-the-loop data annotation approach is illustrated in Figure 8, where the LLM agent is queried to translate Chinese sentences into Nüshu using few-shot prompting. In the initial round (R1), seed translation examples are sampled from the NCGold dataset, while subsequent example sets are drawn from a combination of NCGold and NCSilver after the first round of translations. Beginning in the second round (R2), newly generated samples from the previous round are shuffled, with five samples selected to replace the top 5 samples in the original 35-sample list and the bottom five samples removed. For instance, in R2, the 35 training samples consisted of the top 5 new translations from R1, followed by 30 samples from the original set. By the third round (R3), the 35 samples included five new translations from R2, five from R1, and 25 from the original set. The control test consistently used the original 35 samples across all rounds.
A rule-based translation length validator is integrated into the pipeline to ensure that each Nüshu character maps to a single Chinese character. This validator ensures that the original and translated contents have the same length. If a length mismatch occurs, the generation is rejected, and the LLM is prompted again to resolve the issue. The model is allowed up to 7 retries before moving on and classifying the instance as a failed translation. The enforced one-to-one character translations compensates for limited data, by providing additional linguistic instruction for Nüshu. Our experiments without enforced length validation reported a 31.37% accuracy, versus 48.69% with length validation, showing that data scarcity can be somewhat compensated through incorporating language-specific information.
Figure 8: Pipeline for LLM-in-the-loop data annotation approach (Rounds 1-6).
Figure 9: Translation of unseen characters (boxed in pink).
Figure 10: Translation success and failure across rounds.
@@ 4.2.3 Data Annotation Outcomes
Using only 35 sentences from NCGold as seed examples, the GPT-4-Turbo model achieved an accuracy of 48.69% in perfectly translating the 50 left-out NCGold examples on a strict character-by-character match. A manual review of the translations revealed that minor translation errors were predominantly due to ambiguity in Nüshu character selection, as each Nüshu character can correspond to multiple Chinese characters and vice versa.
This complexity in mapping contributed to most of the discrepancies observed. For out-of-domain Chinese sentences, the GPT-4-Turbo model successfully translated 70.00% of sentences with lengths below 16 characters. However, it frequently refused to respond when the sentence lengths exceeded 31 characters. Figure 10 displays the number of successfully translated sentences and failed translations for each round, along with the results from the control group, which is always provided the original 35 learning samples. Figure 11 exemplifies the correct, incorrect, and untranslatable responses from ChatGPT in our experiments.
@@
Table 2: Evaluation results (averaged) for our Seq2Seq MT Model, evaluated on a fixed 100 test sample.
4.2.4 Other Low-Resource Languages
To evaluate the scalability of our framework, we extended our investigation to additional low-resource languages with somewhat larger datasets than Nüshu but still underrepresented in mainstream language models like GPT-4-Turbo. Cherokee was selected as a representative Native American language, recognized for its moderate usage yet classified as low-resource Zhang et al. (2022). Despite GPT-4-Turbo’s partial familiarity with Cherokee Shu et al. (2024), it struggled with accurate translations. Implementing our framework on a 14k Cherokee-English corpus Zhang et al. (2020) yielded translation accuracies of 28%, 40%, and 42% with 35, 150, and 300 sample sizes, respectively. These results highlight the persistent challenge of achieving higher accuracy in low-resource language translation without tailored linguistic adaptations.
@@ 5 Applications
Leveraging the NCGold and NCSilver datasets, we present both a Nüshu-Chinese language model and a Nüshu-to-Chinese translation model to demonstrate the potential of our data-annotation framework in advancing future research on Nüshu. Our experiments also underscore the critical need for further expanding the Nüshu corpus, enabling researchers to gain deeper insights into the language and the rich cultural heritage it embodies.
@@ 5.1 Language Modeling in Nüshu
Having obtained a set of 598 Nüshu-Chinese translation sentence pairs, we employed the FastText library Athiwaratkun et al. (2018) to train a bilingual skip-gram language model Mikolov et al. (2013) in both Nüshu and Chinese. Specifically, we separated the Nüshu-Chinese translation pairs from NCGold and NCSilver into individual sentences, shuffled them, and utilized this bilingual corpus to train the FastText model.
For the FastText model, we set the embedding vector size to 300, a window size of 10, and a minimum count of 5. We used the skip-gram model (sg=1), which is well-suited for smaller datasets and rare words, as it predicts surrounding words given a target word—particularly useful for the infrequent characters in the Nüshu corpus.
The model was trained for 20 epochs to ensure convergence without over-fitting, with a negative sampling rate of 10 to speed up training by focusing on relevant examples. The default learning rate of 0.05 is used for training our model.
@@ 5.2 Nüshu-to-Chinese Translation
We initially trained a Seq2Seq model for Nüshu-Chinese translation using training samples from NCGold, with a batch size of 3 and for 15 epochs. For the test set, we first shuffled the 500-sentence NCGold and fixed the first 100 samples as the test set. Despite the small dataset size, the model’s performance was promising, comparable to models developed for other truly low-resource languages Baruah et al. (2021).
Encouraged by these results, we combined NCGold with NCSilver, resulting in a total of 598 sentences, to further improve translation quality. While ChatGPT demonstrated potential in translating Chinese text to Nüshu for NCSilver, it remains a closed-source model with limited interpretability, and its outputs are often noisy, requiring multiple queries for acceptable translations. To mitigate these issues and enhance both translation accuracy and model transparency, we retrained the Seq2Seq model using the combined NCGold and NCSilver datasets. We provide further details on this process in Section 6.
@@ 5.3 Nüshu Translates to Other Languages
Given the limited availability of official Nüshu-Chinese translations, data for translating Nüshu into other languages is even more scarce. As mentioned in Section 2, Chiu (2012) published a collection of works featuring Nüshu calligraphy alongside English translations of their corresponding Chinese poems. However, this is not a direct Nüshu-to-English translation and lacks certain linguistic nuances or fidelity Rosa et al. (2017). Nevertheless, it introduces the right concept of using Nüshu as a bridge for multilingual translations.
Our framework enables the construction of larger Nüshu datasets than ever before, making it possible to scale up Nüshu-to-other-language translation datasets when combined with such works. This approach opens up new avenues for expanding Nüshu’s reach across languages and enhancing the diversity of translation datasets.
@@ 6 Discussion
However, we observed that performance peaked at a training dataset of 487 sentences. This dataset includes 400 original sentences from our book translations and 87 sentences generated through Round 5 (R5). Interestingly, performance declined when data from Round 6 (R6) was added. We hypothesize that this decline is due to the sentences in R6 being significantly longer than those in the rest of the dataset. As shown in Table 1, the average sentence length for R6 is 31.73, compared to 23.1 for R5, while the maximum sentence length in the original 500-sentence dataset is only 18.
We hypothesize that the model struggles with excessively long sentences because it was exposed to primarily shorter training data. Seq2Seq models, particularly those using a fixed-length encoder-decoder architecture, tend to perform well on sentence structures and lengths similar to those seen during training. When presented with longer, more complex sentences, the model may struggle to effectively encode the entire sentence into a fixed representation, leading to loss of important information, truncation, or difficulty decoding the translation. Furthermore, attention mechanisms can become less effective with longer sequences, especially when the model has not been fine-tuned or adequately trained. Consequently, incorporating longer sentences without balancing the model’s exposure to various sentence lengths likely contributed to the performance drop observed after R6.
@@ 7 Conclusion and Future Work
In this paper, we introduced NüshuRescue, a novel approach leveraging LLMs to expand text corpora in endangered languages, with Nüshu as a case study. Our method facilitates research on these languages while minimizing reliance on extensive human annotations. Findings show that NüshuRescue can generate high-quality translations from Chinese to Nüshu with only 35 seed pairs, demonstrating its effectiveness in producing accurate translations with minimal input. The model also generalizes well across different sentence lengths and text domains. Notably, NüshuRescue does not require the underlying LLM to have prior knowledge of the target language, making it adaptable to other low-resource languages.
In addition, we digitized 500 Nüshu-Chinese sentence pairs from ancient Chinese poetry, which were used to train both a FastText language model and a Seq2Seq Nüshu-to-Chinese MT model, in combination with the LLM-generated translations. This dataset, encompassing 80% of the only expert-validated Nüshu resource available, is a vital step in preserving Nüshu from extinction. In the absence of other digital resources, our paper serves as a foundational contribution, introducing Nüshu to the NLP community for the first time and providing a data augmentation framework. Looking ahead, NüshuRescue has potential for adaptation for other under-resourced languages, fostering the expansion of corpora and enabling the development of larger, more sophisticated language models.
@@
In this paper, we introduced NüshuRescue, a LLM-based framework designed to scale up text corpora in resource-scarce languages. While our experiments demonstrated its effectiveness and generalizability, we encountered limitations in significantly expanding the Nüshu corpus due to the restricted scope of the current Nüshu-Chinese dictionary. The only available expert-validated source of Nüshu content comes from “A Compendium of Chinese Nüshu” Zhao (1992), which contains only 374 Nüshu characters and 1453 corresponding Chinese characters.
Although NüshuRescue demonstrated a strong ability to generalize to out-of-vocabulary characters, the absence of expert validation [^5] Nüshu no longer has any native speakers, as the last known native speaker passed in 2004. prevents us from safely releasing the generated data. Consequently, we sampled only 180 sentences from the UD Chinese GSDSimp dataset Wang et al. (2021) dataset to construct the NCSilver dataset, ensuring that all included Chinese characters were within the existing Nüshu vocabulary. However, as future studies of Nüshu incorporate more characters, NüshuRescue can be easily adapted to further expand the Nüshu corpus.
@@
In this paper, we exclusively utilized publicly available resources, including the “A Compendium of Chinese Nüshu” book Zhao (1992) and the UD Chinese GSDSimp dataset Wang et al. (2021) for our experiments. We relied on off-the-shelf LLMs, such as GPT-4-Turbo, as the backbone model for NüshuRescue. Furthermore, we have thoroughly documented the prompts used to generate Nüshu translations from the LLMs, detailing how these prompts were iteratively refined to improve the quality of the outputs.
Given these considerations, we do not foresee any ethical concerns arising from our work. Our NCGold and NCSilver datasets, as well as related code, have been made publicly available through our GitHub (https://github.com/ivoryayang/NushuRescue), ensuring transparency and fostering further research in this area.
Acknowledgments
This research was funded in part by the Dartmouth Alumni Research Award. We extend our gratitude to the Dartmouth graduate alumni for their generous support and commitment to fostering academic inquiry.
@@ Appendix A Appendix Figures
Figure A1: An example from the book (a public domain page scan included in the product description on Amazon, and therefore not subject to copyright restrictions).